ng Tol d redctant in solution One way the U.S. Environmental Protection Agency (EPA) tests for chloride contaminants in water is by titrating a sample of silver nitrate solution. Any chloride anions in sol ution will combine with the silver cations to produce bright white silver chloride precipitate. Suppose an EPA chemist tests a 200. mL sample of groundwater known to be contaminated with iront(t) chloride, which would react with sliver nitrate solution like this: FeCl 2(aq) + 2 AgNO3(aa) → 2 AgCI (s) + Fe (NO3)2(aq) The chemist adds 13.0 mM silver nitrate solution to the sample until siver chloride stops forming. She then washes, dries, and weighs the precipitate. She finds she has collected 6.5 mg of silver chloride. Calculate the concentration of iron(I1) chloride contaminant in the original groundwater sample. Be sure your answer has the correct number of significant digits. oP ing
Answer #1

FeCl2 + 2 AgNO3 -----> 2 AgCl + Fe(NO3)2

mass of AgCl obtained = 6.5 mg

number of moles of AgCl = 6.5 * 10^-3 /143.32 = 4.5353 * 10^-5

2 moles of AgCl requires 1 mole of Fecl2

4.5353 * 10^-5 moles of AgCl requires 4.5353 * 10^-5 * 1/2 moles of FeCl2

                                     = 2.2676 * 10^-5 moles of FeCl2

concentration of FeCl2 = 2.2676 * 10^-5/(200*10^-3)

Concentration of FeCl2 = 1.1338 * 10^-4 M = 0.11 mM (to correct number of significant figures)